# Analysis and Simulation of a Fluid-Heat System in a Thin, Rough Layer in Contact With a Solid Bulk Domain

### v2

@@ 5 Simulations
In addition to the previous analysis, we investigate the properties of the effective model with the help of numerical simulations and examine what aspects can be captured in the homogenized problem. In particular, we are looking at the effects of the roughness patterns, viscosity function and inflow conditions. We also study the accuracy of the homogenization limits by direct comparison with the $\varepsilon$-model for small values of $\varepsilon$. Note that we focus on the two-dimensional case for this comparison as the computational cost for the $\varepsilon$-problem is prohibitively high in the case $d=3$.
All simulations are carried out with the finite element solver FEniCS [7] and meshes are created with Gmsh [31]. To stabilize the convection term in our model, a SUPG scheme [15] is applied. We use continuous, piecewise linear finite elements to represent temperature fields $(\theta^{f},\theta^{s})$ and pressure $p$, while the velocity $u$ is modeled with continuous piecewise quadratic elements. The time stepping is performed with a semi-implicit Euler scheme, where we linearized the convection term by using the velocity field from the previous time step. Our implementation is freely available online [24].
We start by collecting all the parameters that are fixed in the following simulations: We consider the time interval $S=(0,5)$ and the macroscopicdomain is the unit square $\Omega=(0,1)^{2}$. The conductivities are set to $\kappa^{s}=0.5$ and $\kappa^{f}=0.1$, and the heat exchange parameter to $\alpha=1.0$. For the initial conditions, everything is set to zero. At $\Sigma$, the Dirichlet condition for the velocity is $u_{\text{motion}}=(1,0)^{T}$. The heat source is given by
$f_{\varepsilon}(x)=\frac{1-\nicefrac{{x_{d}}}{{\varepsilon}}}{1-\gamma_{0}},$
so that there is a higher production at the tip of the roughness, which is motivated by the underlying application of grinding. The temperature-dependent viscosity is modeled by a Vogel-Fulcher-Tammann equation:
@@
which is a reasonable approach for fluids [30]. To obtain a viscosity profile that varies in our temperature range, we choose $\mu_{0}=0.2$, $\alpha=3.0$, and $T_{0}=0.6$. Unless stated otherwise, the time step $\Delta t=0.05$ and the mesh resolution $h=0.03$ are used, where the mesh is locally refined around the thin layer with resolution $h\varepsilon$. A convergence study was carried out with respect to $\Delta t$ and $h$ and the expected linear or quadratic convergence was observed. To study the influence of different roughness functions, we investigated two different cases:
@@
    $\gamma_{\text{rect}}(x_{1};\gamma_{0})=\begin{cases}1\quad&\text{, if }x<0.1\text{ or }x>0.9,\\ 1-(1-\gamma_{0})\left(-2\left(\frac{x_{1}-0.1}{0.1}\right)^{3}+3\left(\frac{x_{1}-0.1}{0.1}\right)^{2}\right)\quad&\text{, if }x\in[0.1,0.2),\\ \gamma_{0}\quad&\text{, if }x\in[0.2,0.8],\\ \gamma_{0}+(1-\gamma_{0})\left(-2\left(\frac{x_{1}-0.8}{0.1}\right)^{3}+3\left(\frac{x_{1}-0.8}{0.1}\right)^{2}\right)\quad&\text{, if }x\in(0.8,0.9].\end{cases}$
@@
Table 1: Roughness patterns used in the numerical simulations, and the corresponding effective parameters. Since the setup is two-dimensional, all parameters are reduced to scalar values.
Three different inflow patterns are compared, a linear inflow, a quadratic, and a linear inflow that is not zero above $\gamma_{0}$ given by
$(i)\ \ u_{\text{BC, lin}}=\begin{cases}\left(1-\frac{x_{2}}{\gamma_{0}}\right)u_{\text{motion}}&\text{, if }x_{2}\leq\gamma_{0},\\ 0&\text{, if }x_{2}>\gamma_{0},\end{cases}\quad$ $,\quad(ii)\ \ u_{\text{BC, quad}}=\begin{cases}\left(1-\left(\frac{x_{2}}{\gamma_{0}}\right)^{2}\right)u_{\text{motion}}&\text{, if }x_{2}\leq\gamma_{0},\\ 0&\text{, if }x_{2}>\gamma_{0},\end{cases}$
and ($\text{and }(iii)\ \ u_{\text{BC, lin, 2}}$ $=(1-x_{2})u_{\text{motion}}.$
The effective inflow then equals $\bar{u}_{\text{BC, lin}}=\frac{1}{2}\gamma_{0}$, $\bar{u}_{\text{BC, quad}}=\frac{2}{3}\gamma_{0}$ or $\bar{u}_{\text{BC, lin, 2}}=\frac{1}{2}$. With $u_{\text{BC, lin, 2}}$ we investigate how the model behaves if Assumption Item (A5) is not fulfilled. In this case, we can not build an extension of the boundary condition to the whole domain by a simple scaling argument. Instead, the extension has to be carried out in the thin rough layer directly. If not specified otherwise, we use $u_{\text{BC, lin}}$ in the following simulation studies.
@@
Please note that the overall simulation scenarios are directly motivated by the grinding application in the way we chose the boundary conditions (inflow/outflow), the roughness patterns, and the viscosity functions. However, these simulation experiments are only qualitative, as we are not working with realistic and sensible parameter values for the conductivities, the heat source, etc. We refer to [57] where quantitative simulations using realistic material parameters were performed for a single grain scenario.
Comparison between $\varepsilon$-model and effective model. We start by comparing the solutions to the effective model with solutions of the resolved $\varepsilon$-model (where we generally compare with values $\varepsilon=0.1,0.05.,0.01$) for the sinusoidal roughness pattern and $\gamma_{0}=0.5$. To give a general overview of the solution, we present them in Fig. 2 and Fig. 3(d). In the following, we focus mainly on comparing the solutions inside the thin layer as this is the critical area where the interesting features (rough boundary and thin domain) are present. As the solutions are defined on different domains (the effective solutions are defined on a line and the $\varepsilon$-solutions on a thin domain), we take an average of the microscopic solution with respect to the vertical direction and compare this with the corresponding effective quantities, as done in similar numerical studies [40, Section 5]. To this end, we define the averaged functions
$\begin{split}\bar{\theta}^{f}_{\varepsilon}(t,\tilde{x})\coloneqq\frac{1}{\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}&\int_{0}^{\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}\theta^{f}_{\varepsilon}(t,\tilde{x},s)\,\mathrm{d}s,\quad\bar{p}_{\varepsilon}(t,\tilde{x})\coloneqq\frac{\varepsilon}{\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}\int_{0}^{\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}p_{\varepsilon}(t,\tilde{x},s)\,\mathrm{d}s\\ &\bar{u}_{\varepsilon}(t,\tilde{x})\coloneqq\frac{1}{\varepsilon}\int_{0}^{\varepsilon\gamma(\nicefrac{{\tilde{x}}}{{\varepsilon}})}u_{\varepsilon}(t,\tilde{x},s)\,\mathrm{d}s.\end{split}$ (5.1)
Note that $\bar{p}_{\varepsilon}$ is scaled with $\varepsilon^{2}$ ($\nicefrac{{\varepsilon^{2}}}{{\varepsilon\gamma}}=\nicefrac{{\varepsilon}}{{\gamma}}$) since the effective pressure $p$ is the limit of $\varepsilon^{2}p_{\varepsilon}$. We observe that the effective model captures the resolved $\varepsilon$-model fairly well with regard to both temperature and pressure, as observable in Fig. 2 to 4. We obtain negative pressure at the inflow ($x_{1}=0$) since the amount of fluid pushed into the domain by $u_{BC,lin}$ is less than the fluid movement induced by the bottom boundary. This becomes also apparent, by comparing $\bar{u}_{\text{BC, lin}}$ with the flow induced by the cell problem (4.2), which gives the value $\bar{\xi}_{0}=0.29$ (see Table 1).
Note that for velocity, the average of the microscopic solution is already close to the effective value of $\bar{u}_{\text{BC, lin}}=\frac{1}{2}\gamma_{0}=0.25$ and only oscillates slightly around this value, see Fig. 4. These oscillations originate mainly from numerical inaccuracies, induced both by the FE-scheme and the numeric integration to compute Eq. 5.1.
@@
Figure 2: Comparison of the temperature profiles, at $t=5$, for the micro and effective model, with sinusoidal roughness. The effective fluid temperature $\theta^{f}$ is only defined on the line $(0,1)\times\{0\}$ but for visualization purposes the line is being shown with a thickness
@@
Figure 3: Comparison of the fluid (top) and pressure (bottom) profiles, at $t=5$, for the micro and effective model, with sinusoidal roughness. The line in the effective case is again visualized with a thickness
@@
Figure 4: Comparison inside the thin layer of the effective solutions with the averaged micro solutions defined in Eq. 5.1. All solutions are shown at the final time step $t=5$ and plotted over $\Sigma$. The velocity field for $\varepsilon=0.01$ is left out for a clearer visualization.
We also want to study numerically the convergence behavior for $\varepsilon\to 0$. To this end, we compute inside the thin layer the difference of the effective and averaged solution in the $L^{2}(S\times\Sigma)$ norm, and in the solid bulk domain, we compare in the $L^{2}(S\times\Omega^{s}_{\varepsilon})$ norm. The results are shown in Fig. 5. For temperature and pressure, we observe a convergence of order $\mathcal{O}(\varepsilon)$. On the other hand, there is no observable convergence for the velocity, as mentioned before, we assume that this is mainly due to numerical errors. The difference between effective and averaged velocity is also already small compared to temperature and pressure error.
Figure 5: Convergence study for the limit $\varepsilon\to 0$. Both axes are on a logarithmic scale. Inside the thin layer, meaning for $\theta^{f},p,u$, we present the error in the $L^{2}(S\times\Sigma)$ norm. The error for $\theta^{s}$ is computed in $L^{2}(S\times\Omega^{s}_{\varepsilon})$. To estimate the order of convergence, the black line equals the identity.
@@
Varying the inflow function. To study how well the effective model captures the influence of the prescribed inflow, we compare different Dirichlet conditions for the roughness given by $\gamma_{\text{sine}}$ for $\gamma_{0}=0.5$ and fixed $\varepsilon=0.05$. The resulting temperature, pressure, and velocity profiles at time $t=5$ are visualized in Fig. 6. We first observe that, naturally, a higher average fluid flow leads to a cooler system since the heat energy is removed faster from the system. Furthermore, inflows $u_{\text{BC, quad}}$ and $u_{\text{BC, lin, 2}}$ result in a positive pressure at the inflow interface, since the amount of fluid pressed into the domain is larger than the amount moved by the bottom boundary condition. The homogenized model can capture the influence of the different inflow conditions fairly well with regard to temperature and pressure. This holds true even for $u_{\text{BC, lin, 2}}$, which does not satisfy the assumptions made in Section 2.3.
@@
Figure 6: Comparison of the fluid temperature, pressure, and horizontal velocity component for different inflow functions, at $t=5$. In black the micro solution, and in orange the effective results. The velocity is visualized along $\{0.5\}\times(0,\varepsilon)$. For the other functions, we again compare the average (5.1).
@@
Varying the grain height $\gamma_{0}$. We investigate the influence of the minimal height of the interface $\gamma_{0}$. Here, we again focus on the roughness given by $\gamma_{\text{sine}}$ and fixed $\varepsilon=0.05$. We use the inflow function $u_{\text{BC, lin, 2}}$ as it is independent of $\gamma_{0}$; this allows us to focus on the influence of the interface height. Similarly, we modify the heat source to $f_{\varepsilon}=\tfrac{1}{|\Gamma|}$ ensuring that the total amount of heat produced is one and is independent of $\gamma_{0}$. The simulated solutions, again at $t=5$, are presented in Fig. 7. As in the results before, the homogenized model shows good agreement with the exact solutions for $\varepsilon=0.05$. The grain height has a strong influence on the pressure profile; since we have a fixed amount of inflow for all cases, a smaller gap necessarily leads to higher pressures inside the layer. At the same time, the fluid velocity in between the ”humps” created by the roughness is also reduced, leading to slightly higher temperatures.
@@
Varying the roughness pattern. In this simulation scenario for the two-dimensional case, we demonstrate the influence of different roughness patterns. We utilize the same setup as in the previous case for varying roughness height but this time we fix $\gamma_{0}=0.5$ and compare the differences between $\gamma_{\text{sine}}$ and $\gamma_{\text{rect}}$. As before, we present the results at the last time step, in Fig. 8. The homogenized model accurately reflects the influence of the roughness pattern on all solution fields. For 2D we can conclude, that the derived effective model yields a good approximation of the resolved micro model and the temperature-dependent viscosity has a, rather large, impact on the pressure profile.
@@
$f_{\varepsilon}(x)=\begin{cases}1\quad&\text{, if }x_{2}>0.5\text{ and }|x_{1}-0.5|<0.1,\\ 0&\text{ else}.\end{cases}$
This right-hand side is chosen in such a way that local variations in the heat production occur, leading to local differences in the viscosity. The inflow is defined by $u_{\text{BC, lin}}$ with $\gamma=0.5$ and $u_{\text{motion}}=(1,0,0)^{T}$. For the roughness we define

### v3

@@ 5 Simulations
In addition to the previous analysis, we investigate the properties of the effective model with the help of numerical simulations and examine what aspects can be captured in the homogenized problem. In particular, we are looking at the effects of the roughness patterns, viscosity function, and inflow conditions. We also study the accuracy of the homogenization limit by direct comparison with the $\varepsilon$-model for small values of $\varepsilon$. Note that we focus on the two-dimensional case for this comparison as the computational cost for the $\varepsilon$-problem is prohibitively high in the case $d=3$.
All simulations are carried out with the finite element solver FEniCS [53] and meshes are created with Gmsh [54]. To stabilize the convection term in our model, a SUPG scheme [55] is applied. We use continuous, piecewise linear finite elements to represent temperature fields $(\theta^{f},\theta^{s})$ and pressure $p$, while the velocity $u$ is modeled with continuous piecewise quadratic elements. The time stepping is performed with a semi-implicit Euler scheme, where we linearize the convection term by using the velocity field from the previous time step. Our implementation is freely available online [39].
We start by collecting all the parameters that are fixed in the following simulations: We consider the time interval $S=(0,5)$, and the macroscopic domain is the unit square $\Omega=(0,1)^{2}$. The conductivities are set to $\kappa^{s}=0.5$ and $\kappa^{f}=0.1$, and the heat exchange parameter to $\alpha=1.0$. The initial conditions are set to zero. At $\Sigma$, the Dirichlet condition of the velocity is set to $u_{\text{motion}}=(1,0)$. The heat sources are given by $f^{f}_{\varepsilon}=0$ and
$f_{\varepsilon}^{s}(x)=\frac{1-\nicefrac{{x_{d}}}{{\varepsilon}}}{1-\gamma_{0}},$
so that there is a higher production at the tip of the roughness, which is motivated by the underlying application of grinding and the fluid only heats up through exchange with the Robin condition. The temperature-dependent viscosity is modeled by a Vogel-Fulcher-Tammann equation:
@@
which is a reasonable approach for fluids [5]. To obtain a viscosity profile that varies in our temperature range, we choose $\mu_{0}=0.2$, $\alpha=3.0$, and $T_{0}=0.6$. Unless stated otherwise, the time step $\Delta t=0.05$ and the mesh resolution $h=0.03$ are used, where the mesh is locally refined around the thin layer with resolution $h\varepsilon$. A convergence study was carried out with respect to $\Delta t$ and $h$ and the expected linear or quadratic convergence was observed. To study the influence of different roughness patterns, we model the solid-fluid interface by the graph of a periodic function $\gamma:Y^{d-1}\to[\gamma_{0},1]$, meaning $\Gamma=\{(\tilde{x},\gamma(\tilde{x})):\tilde{x}\in Y^{d-1}\}$. We investigate two different setups:
@@
[…]
@@
Table 1: Roughness patterns used in the numerical simulations, and the corresponding effective parameters. Since the setup is two-dimensional, all effective parameters reduce to scalar values.
@@
Three different inflow patterns are compared: a linear inflow, a quadratic inflow, and a linear inflow that is not zero above $\gamma_{0}$ given by
$(i)\ \ u_{\text{BC, lin}}=\begin{cases}\left(1-\frac{x_{2}}{\gamma_{0}}\right)u_{\text{motion}}&\text{, if }x_{2}\leq\gamma_{0},\\ 0&\text{, if }x_{2}>\gamma_{0},\end{cases}$
$(ii)\ \ u_{\text{BC, quad}}=\begin{cases}\left(1-\left(\frac{x_{2}}{\gamma_{0}}\right)^{2}\right)u_{\text{motion}}&\text{, if }x_{2}\leq\gamma_{0},\\ 0&\text{, if }x_{2}>\gamma_{0},\end{cases}$
$(iii)\ \ u_{\text{BC, lin, 2}}=(1-x_{2})u_{\text{motion}}.$
The effective inflow then equals $\bar{u}_{\text{BC, lin}}=\frac{1}{2}\gamma_{0}$, $\bar{u}_{\text{BC, quad}}=\frac{2}{3}\gamma_{0}$ or $\bar{u}_{\text{BC, lin, 2}}=\frac{1}{2}$. With $u_{\text{BC, lin, 2}}$ we construct an inflow that does not fulfill (A5) (the inflow is nonzero also for $x_{2}>\gamma_{0}$). Hence, we cannot use a simple scaling argument to build an extension of the boundary condition to the whole domain. However, we observe that the effective model also works well for this larger class of functions. If not specified otherwise, we use $u_{\text{BC, lin}}$ in the following simulation studies.
@@
Please note that the overall simulation scenarios are motivated by the grinding application. However, these simulation experiments are only qualitative, as we are not working with realistic and sensible parameter values for the conductivities, the heat source, etc., and instead only highlight the behavior of the effective model. We refer to [8] where quantitative simulations using realistic material parameters were performed for a single-grain scenario.
Comparison between $\varepsilon$-model and effective model. We start by comparing the solutions to the effective model with solutions of the resolved $\varepsilon$-model (where we compare with $\varepsilon=0.2,0.1,0.05,0.01$) for the sinusoidal roughness pattern and $\gamma_{0}=0.5$. We present a general overview of the solution in Fig. 2 and Fig. 3. Note that we multiply the microscale pressure $p_{\varepsilon}$ with $\varepsilon^{2}$, since convergence only holds for $\varepsilon^{2}p_{\varepsilon}$. In the following, we focus mainly on comparing the solutions inside the thin layer, as this is the critical area where the interesting features (rough boundary and thin domain) are present.
We observe that the effective model captures the temperature and pressure of the resolved $\varepsilon$-model fairly well (Fig. 2 and 3). To better observe the microscopic fluctuations of temperature and pressure as well as the effective profile of the effective solution, the micro solution is visualized along the line $(0,1)\times\{\nicefrac{{\varepsilon}}{{2}}\}$ in Figure 4 and compared with the effective solution. We obtain negative pressure at the inflow ($x_{1}=0$) since the amount of fluid pushed into the domain by $u_{BC,lin}$ is less than the fluid movement induced by the bottom boundary. This becomes also apparent, by comparing $\bar{u}_{\text{BC, lin}}$ with the flow induced by the cell problem (4.2.2), which gives the value $\bar{\xi}_{0}=0.29$ (see Table 1).
@@
Figure 2: Comparison of the temperature profiles, at $t=5$, for the micro and effective model, with sinusoidal roughness. The effective fluid temperature $\theta^{f}$ is only defined on the line $(0,1)\times\{0\}$, but for visualization purposes, the line is being shown with a thickness.
@@
Figure 3: Comparison of the fluid (top) and pressure (bottom) profiles, at $t=5$, for the micro and effective model, with sinusoidal roughness. The line in the effective case is again visualized with a thickness.
@@
Figure 4: Comparison of the effective solutions with the microscopic temperature and pressure inside the layer. The micro solution is plotted along the line $(0,1)\times\{\nicefrac{{\varepsilon}}{{2}}\}$. All solutions are shown at the final time step $t=5$.
We also want to study numerically the convergence behavior for $\varepsilon\to 0$ explicitly and not only compare visually. The solutions, however, are defined on different domains (the effective solutions are defined on a line and the $\varepsilon$-solutions on a thin domain), but given the results in Section 4, the functions in the thin layer can be approximated by the effective solutions via
[…]
The functions $\theta^{f}_{1}$ and $p_{1}$ are defined in Equation 27 and (39) and can be computed with the cell solutions given by Equation 28, (4.2.2), (4.2.2) and the homogenized solution. With the reconstruction (42), we compute the difference of the micro solution with the effective solution in the $L^{2}(S\times\Omega^{f}_{\varepsilon})$ norm. The results are shown in Fig. 5. For temperature and pressure, we observe a convergence of order $\mathcal{O}(\varepsilon^{2})$. On the other hand, the velocity converges only with order $\mathcal{O}(\varepsilon)$, which is to be expected given that the temperature and pressure reconstruction also include higher order terms compared to the velocity. Additionally, the temperature in the solid $\theta^{s}_{\varepsilon}$ converges also with order $\mathcal{O}(\varepsilon)$ in $L^{2}(S\times\Omega^{s}_{\varepsilon})$, but is not depicted for clearer visualization.
Figure 5: Convergence study for the limit $\varepsilon\to 0$. Both axes are on a logarithmic scale. The error between microscopic resolved simulations and the effective quantities reconstructed with Equation 42 in the $L^{2}(S\times\Omega^{f}_{\varepsilon})$ norm. To better estimate the order of convergence, the black lines show different powers of $\varepsilon$.
@@
Varying the inflow function. To study how well the effective model captures the influence of the prescribed inflow, we compare different Dirichlet conditions for the roughness given by $\gamma_{\text{sine}}$ for $\gamma_{0}=0.5$ and fixed $\varepsilon=0.05$. The resulting temperature, pressure, and velocity profiles at time $t=5$ are visualized in Fig. 6. Here we keep the visualization scheme from Figure 4 for the temperature and pressure to highlight the effective solution profile, and not explicitly compare the error in any norm. For the fluid velocity, we compare along the vertical line $\{0.5\}\times(0,\varepsilon)$ and use Equation 42 to compute the effective profile. We first observe that, naturally, a higher average fluid flow leads to a cooler system since the heat energy is removed faster from the system. Furthermore, inflows $u_{\text{BC, quad}}$ and $u_{\text{BC, lin, 2}}$ result in a positive pressure at the inflow interface, since the amount of fluid pressed into the domain is larger than the amount moved by the bottom boundary condition. The homogenized model can capture the influence of the different inflow conditions fairly well with regard to all quantities. This holds true even for $u_{\text{BC, lin, 2}}$, which does not directly satisfy the assumptions made in (A5).
@@
Figure 6: Comparison of the fluid temperature, pressure, and horizontal velocity component for different inflow functions, at $t=5$. In black, the micro solution, and in orange, the effective results. The velocity is compared along $\{0.5\}\times(0,\varepsilon)$.
@@
Varying the grain height $\mathbf{\gamma_{0}}$. We investigate the influence of the minimal height of the interface $\gamma_{0}$. Here, we again focus on the roughness given by $\gamma_{\text{sine}}$ and fixed $\varepsilon=0.05$. We use the inflow function $u_{\text{BC, lin, 2}}$ as it is independent of $\gamma_{0}$; this allows us to focus on the influence of the interface height. Similarly, we modify the heat source to $f_{\varepsilon}=\tfrac{1}{|\Gamma|}$, ensuring that the total amount of heat produced is one and is independent of $\gamma_{0}$. The simulated solutions, again at $t=5$, are presented in Fig. 7. As in the results before, the homogenized model shows good agreement with the exact solutions for $\varepsilon=0.05$. Given that we plot along the line at $x_{2}=\nicefrac{{\varepsilon}}{{2}}$, there appear jumps for the case $\gamma_{0}=0.1$ since the line is not always inside the domain. The grain height has a strong influence on the pressure profile; since we have a fixed amount of inflow for all cases, a smaller gap necessarily leads to higher pressures inside the layer. At the same time, the fluid velocity in between the ”humps” created by the roughness is also reduced, leading to slightly higher temperatures.
@@
Varying the roughness pattern. In this simulation scenario for the two-dimensional case, we demonstrate the influence of different roughness patterns. We utilize the same setup as in the previous case for varying roughness height, but this time we fix $\gamma_{0}=0.5$ and compare the differences between $\gamma_{\text{sine}}$ and $\gamma_{\text{rect}}$. As before, we present the results at the last time step, in Fig. 8. The homogenized model accurately reflects the influence of the roughness pattern on all solution fields. For the 2D case, we can conclude that the derived effective model yields a good approximation of the resolved micro model, and the temperature-dependent viscosity only has an impact on the pressure profile.
@@
$f_{\varepsilon}^{s}(x)=\begin{cases}1\quad&\text{, if }x_{2}>0.5\text{ and }|x_{1}-0.5|<0.1,\\ 0&\text{ else}.\end{cases}$
This right-hand side is chosen in such a way that local variations in the heat production occur, leading to local differences in the viscosity. The inflow is defined by $u_{\text{BC, lin}}$ with $\gamma_{0}=0.5$ and $u_{\text{motion}}=(1,0,0)^{T}$. For the roughness, we again model the reference cell as a graph and define